# Topology of Discrete Quantum Feedback Control

### v1

@@ IV.3 Topological phase transition
The chiral Maxwell’s demon described in Sec. IV.1 exhibits topological phase transitions as the duration $\tau$ of feedback control is changed. In Fig. 8, we show the eigenspectrum under the PBC for four different values of the duration of feedback control. As the duration of feedback control is increased, the absolute value of the winding number also increases. In Fig. 9(a), we show the dependence of the winding number on the duration of feedback control. The topological phase transitions occur at an almost constant rate as a function of the duration.
@@
Figure 8: Eigenspectrum of the quantum channel for the chiral Maxwell’s demon for four different values of the duration of the feedback. A blue curve and a blue dot constitute the eigenspectrum under the PBC in the limit of infinite system size, and orange dots show the eigenspectrum under the OBC with the system size of $L=50$. The unit of energy is set to $J=1$, and we take $V=+\infty$. The duration of feedback control is (a) $\tau=0.5$, (b) $\tau=2$, (c) $\tau=3$, and (d) $\tau=3.9$. The winding number $w$ around the origin is (a) $w=0$, (b) $w=-2$, (c) $w=-3$, and (d) $w=-4$.
@@
Figure 9: (a) Dependence of the winding number $w(\xi_{\mathrm{PG}}=0)$ on the duration $\tau$ of feedback control. (b) Velocity of a particle during feedback control. The height of the potential is set to be $V=+\infty$.
@@
The topological phase transitions can physically be understood from the ballistic transport of a particle during the feedback control. In Fig. 9(b), we show the mean velocity
@@
of a particle during the feedback control. As seen from the figure, the particle shows ballistic transport at a constant velocity after an initial transient dynamics. To unveil the relation between the ballistic tranport and the topological phase transitions, we rewrite the winding number (46) by using $z=e^{-ik}$ as
@@
is the probability-generating function of $\{p(r)\}$. From Eq. (48) and by using the argument principle, we have
@@
The topological phase transitions are accompanied by exceptional points, at which the CPTP map cannot be diagonalized [68]. In fact, when the value of the winding number around $\xi_{\mathrm{PG}}=0$ changes, $\xi_{0}(k)$ should vanish for some momentum $k$, and the Bloch matrix $X(k)$ cannot be diagonalized at that point (see Appendix B). Thus, here topology causes the emergence of exceptional points of a CPTP map.
@@ V Symmetry classification of feedback control
Symmetry enriches topological phases since two topological phases may be distiguished in the presence of symmetry [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we discuss symmetry of quantum channels for discrete quantum feedback control. In Sec. V.1, we discuss unitary symmetry. In Sec. V.2, we consider an antiunitary symmetry associated with every quantum channel. While the contents in these two subsections can be found in literature [90, 91, 33, 92, 66, 67], we reproduce them to make the present paper self-contained. In Sec. V.3, we discuss the Bernard-LeClair (BL) symmetry classes of quantum channels for discrete quantum feedback control. In particular, we find that only a subset of the BL symmetry classes are compatible with projective measurement. Consequently, we obtain the ten-fold symmetry classification of discrete quantum feedback control, as listed in Table 2.
@@ V.1 Unitary symmetry
We first consider unitary symmetry of quantum channels. Similarly to the case of open quantum systems described by the Lindblad equation [90, 91], a unitary symmetry of a quantum channel can be either weak symmetry or strong symmetry [92]. A CPTP map $\mathcal{E}$ has a unitary symmetry if it commutes with a unitary superoperator $\mathcal{V}$, that is,
@@
Suppose that each Kraus operator of a CPTP map commutes with a unitary operator $\hat{V}$ as $[\hat{K}_{m},\hat{V}]=0\ (\forall m)$. Then, the CPTP map satisfies Eq. (52) with $\tilde{\mathcal{V}}=\hat{V}\otimes\hat{I}$ or $\tilde{\mathcal{V}}=\hat{I}\otimes\hat{V}^{*}$. Such symmetry is called strong symmetry [90, 91, 92], which gives a more stringent condition than weak symmetry. Strong symmetry is related to a conserved quantity of the dynamics. To see this, we consider the case of continuous symmetry and assume that each Kraus operator $\hat{K}_{m}$ commutes with $\hat{V}_{\theta}=\exp[-i\theta\hat{A}]$ for all $\theta\in\mathbb{R}$, where $\hat{A}$ is a Hermitian operator. This assumption implies $[\hat{K}_{m},\hat{A}]=0\ (\forall m)$. Then, for an arbitrary density matrix $\hat{\rho}$, we have
@@ V.2 Modular conjugation symmetry
Namely, the modular conjugation operator $\tilde{J}$ is a combination of a swap operation on states and the complex conjugation of their coefficients. The modular conjugation satisfies $\tilde{J}^{2}=1$ and acts on the vectorized density matrix [see Eq. (22)] as
@@ V.3 Ten-fold symmetry classification of feedback control with projective measurement
The topological classification of non-Hermitian matrices is performed based on the Bernard-LeClair (BL) symmetry classes [75], which are the non-Hermitian generalization [34, 35] of the Altland-Zirnbauer (AZ) symmetry classes [94]. Here, we discuss the BL symmetry classes of quantum channels for discrete quantum feedback control. In general, a symmetry of a CPTP map $\mathcal{E}$ that belongs to a BL symmetry class has one of the following representations [68]:
@@
where $\hat{P}_{m}$ is a projection operator satisfying $\hat{P}_{m}^{\dagger}=\hat{P}_{m}$ and $\hat{P}_{m}\hat{P}_{m^{\prime}}=\delta_{m,m^{\prime}}\hat{P}_{m}$. An important property of the projective measurement channel (59) is that its eigenvalues are either zero or one, which follows from the Hermiticity and idempotence of projection operators (see Appendix C for the proof). Physically, this property is related to the repeatability of projective measurements, that is, repeated projective measurements after a projective measurement with outcome $m$ always yield the same outcome $m$ [47]. To identify the possible symmetry classes of the projective measurement channel (59), we note that a certain BL symmetry places a constraint on the eigenvalues of a CPTP map. Let $\xi$ be an eigenvalue of a CPTP map $\mathcal{E}$. Then, if this CPTP map has $P$ symmetry (58a) or $C$ symmetry with $\epsilon_{C}=-1$, $-\xi$ is also an eigenvalue of $\mathcal{E}$ [68]. Similarly, if a CPTP map $\mathcal{E}$ has $K$ ($Q$) symmetry, $\epsilon_{K}\xi^{*}$ ($\epsilon_{Q}\xi^{*}$) is also an eigenvalue of $\mathcal{E}$. Since a projective measurement channel (59) always has an eigenvalue $+1$ but cannot have an eigenvalue $-1$, we conclude that $P$ symmetry and $X=C,K,Q$ symmetry with $\epsilon_{X}=-1$ cannot be a symmetry of the projective measurement channel (see Fig. 10). Thus, the possible symmetries of a CPTP map for projective measurement are the following ones:
@@
Figure 10: Eigenspectrum of a projective measurement channel (shown by green dots) consisting of eigenvalues zero and one. Since a projective measurement channel cannot have eigenvalue $-1$, its possible BL symmetry classes are restricted to a subset of the entire 38 symmetry classes.
Now let us consider quantum channels for discrete quantum feedback control with projective measurement. Specifically, here we consider a CPTP map of the form
@@
We list the ten-fold symmetry classification of quantum channels in Table 2, where each symmetry class is labeled according to the general classification of non-Hermitian symmetry classes in Ref. [34]. We note that these ten symmetry classes are equivalent to the ${}^{\dagger}$ classes in Ref. [34] if we multiply $i$ to the original CPTP maps. In fact, the $C$ symmetry (60a) and the $K$ symmetry (60b) are equivalent to “time-reversal ${}^{\dagger}$” and “particle-hole ${}^{\dagger}$” of $i\tilde{\mathcal{E}}$, respectively [95, 34]. In Table 2, we also show the ${}^{\dagger}$ classes corresponding to the symmetry classes of CPTP maps.
Once the relevant symmetry classes are identified, we can exploit the topological classification of general non-Hermitian matrices [34, 35] to examine whether there exists topological feedback control that is described by a topologically nontrivial Bloch matrix in a given symmetry class and spatial dimension. We note, however, that the Bloch matrix cannot be a general non-Hermitian matrix due to the completely positivity and trace-preserving nature of quantum channels (see Appendix A.4). In particular, the complete positivity may place nontrivial constraints on possible topological phases realized with feedback control. We will present an example of this phenomenon in Sec. VI.2.
The present symmetry classification applies to the case with multipoint projective measurements. For example, if a CPTP map for feedback control with a two-point projective measurement is written as
@@
Even if a measurement is not projective, other symmetry classes are still not allowed as long as eigenvalues of a CPTP map for measurement do not always appear as pairs like $(\xi,-\xi)$ or $(\xi,-\xi^{*})$. For example, if a measurement is not projective due to an error in the measurement process, the possible symmetry classes are not altered because we can take an error-free limit under which the symmetry is intact.
The ten ${}^{\dagger}$ symmetry classes are formally equivalent to the symmetry classes for generators of Markovian quantum dynamics discussed in Ref. [65]. In fact, the symmetry classes that are given in Table 2 are consistent with the symmetry classes of CPTP maps that are generated by Lindbladians (see Appendix D). However, we note that a CPTP map for feedback control cannot have a generator if it has a zero eigenvalue [see Eq. (39)]. In general, a CPTP map generated by a Lindbladian can be classified by the 38 BL symmetry classes if one includes shifted symmetries [96, 66, 67]. However, a projective measurement channel (59) cannot have those shifted symmetries because the number of zero eigenvalues is different from that of unit eigenvalues, while shifted symmetry is accompanied by the symmetry of the eigenspectrum with respect to a certain line [96, 66, 67]. Thus, we conclude that the ten-fold symmetry classification exhausts all relevant symmetry classes of discrete quantum feedback control with projective measurement.
We note that a CPTP map always has the modular conjugation symmetry [see Eq. (57) in Sec. V.2]. The modular conjugation symmetry corresponds to a $K$ symmetry with $\eta_{K}=+1$. Thus, for symmetry classes that do not include $K$ symmetry with $\eta_{K}=+1$, we require that a CPTP map has a unitary symmetry and consider the BL symmetry class of each symmetry sector, which does not necessarily have the modular conjugation symmetry. We will revisit this point after giving concrete examples in Sec. VI.
Here, we discuss sufficient conditions for the BL symmetries of CPTP maps to satisfy Eqs. (60a)-(60c), which are helpful to gain an insight into feedback control with symmetry. The symmetries in Eqs. (60a), (60b), and (60c) are satisfied if Kraus operators satisfy

### v2

@@ IV.3 Topological transition
The chiral Maxwell’s demon described in Sec. IV.1 exhibits topological transitions as the duration $\tau$ of feedback control is changed. In Fig. 8, we show the eigenspectrum under the PBC for four different values of the duration of feedback control. As the duration of feedback control is increased, the absolute value of the winding number also increases. In Fig. 9(a), we show the dependence of the winding number on the duration of feedback control. The topological transitions occur at an almost constant rate as a function of the duration of feedback control.
@@
Figure 8: Eigenspectrum of the quantum channel for the chiral Maxwell’s demon for four different values of the duration of the feedback. A blue curve and a blue dot constitute the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for the system size of $L=50$. Energy is measured in units of $J=1$, and the height of the potential is set to $V=+\infty$. The duration of feedback control is (a) $\tau=0.5$, (b) $\tau=2$, (c) $\tau=3$, and (d) $\tau=3.9$. The winding number $w$ around the origin is (a) $w=0$, (b) $w=-2$, (c) $w=-3$, and (d) $w=-4$.
@@
Figure 9: (a) Dependence of the winding number $w(\xi_{\mathrm{PG}}=0)$ on the duration $\tau$ of feedback control. (b) Velocity of a particle during feedback control [see Eq. (47)]. The height of the potential is set to be $V=+\infty$.
@@
The topological transitions can physically be understood from the ballistic transport of a particle during the feedback control. In Fig. 9(b), we show the mean velocity
@@
of a particle during the feedback control. As seen from the figure, the particle shows ballistic transport at a constant velocity after an initial transient dynamics. To understand the relation between the ballistic tranport and the topological transitions, we rewrite the winding number (46) by using $z=e^{-ik}$ as
@@
is the probability-generating function of $\{p(r)\}$, and the integration countor in Eq. (48) is taken in the counterclockwise direction. From Eq. (48) and by using the argument principle, we have
@@
We note that the mean displacement
$\bar{r}:=$ $\sum_{j}\sum_{r}rp(j+r|j)p_{j}$
$=$ $\sum_{r}rp(r)$
[…]
during the topological feedback control is not quantized in general, as inferred from Fig. 9(b). Thus, a quantized topological invariant of a quantum channel does not necessarily guarantee the quantization of a physical observable. This situation is analogous to that of other topological nonequilibrium dynamics [21, 30] such as nonadiabatic topological pumping [22].
The topological transitions are accompanied by exceptional points, at which the CPTP map cannot be diagonalized [68]. In fact, when the value of the winding number around $\xi_{\mathrm{PG}}=0$ changes, $\xi_{0}(k)$ should vanish for some momentum $k$, and the Bloch matrix $X(k)$ cannot be diagonalized at that point (see Appendix B). Thus, topology causes the emergence of exceptional points of a CPTP map.
IV.4 Robustness of chiral transport against disorder
The unidirectional transport induced by the chiral Maxwell’s demon is of a topological origin and therefore robust against external perturbations. Here we demonstrate that the topological feedback control enhances the stability of particle transport. By way of illustration, we consider a task of transporting a paricle from the leftmost site to the rightmost site in an $L$-site chain by using the chiral Maxwell’s demon. Suppose that the feedback Hamiltonian $\hat{H}_{j}$ is perturbed by onsite disorder potentials as
[…]
$+\sum_{i=1}^{L}W_{i}\ket{i}\bra{i},$ (52)
where $W_{i}\in\mathbb{R}$ denotes a random potential. For comparison, we consider a non-topological feedback control with a quantum channel
$\mathcal{E}^{\prime}(\hat{\rho})=\sum_{m=0,1}e^{-i\hat{H}_{m}^{\prime}\tau}\hat{M}_{m}^{\prime}\hat{\rho}\hat{M}_{m}^{\prime{\dagger}}e^{i\hat{H}_{m}^{\prime}\tau},$ (53)
where measurement operators are given by $\hat{M}_{0}^{\prime}=\hat{I}-\ket{L}\bra{L}$ | and $\hat{M}_{1}^{\prime}=\ket{L}\bra{L}$ | and the feedback Hamiltonians are given by
$\hat{H}_{0}^{\prime}=$ $|\displaystyle-J\sum_{i=1}^{L-1}(\ket{i}\bra{i+1}+\mathrm{H.c.})+\sum_{i=1}^{L}W_{i}\ket{i}\bra{i}$ | (54a)
and
$\hat{H}_{1}^{\prime}=$ $0,$ (54b)
i.e., only the rightmost site is measured projectively and the dynamics is stopped when a particle is found at that site.
In Fig. 10, we show the dynamics of the expected value of the position
$\langle\hat{x}\rangle_{N}:=\mathrm{Tr}[\hat{x}\mathcal{E}^{N}(\hat{\rho_{\mathrm{i}}})]$ (55)
with $\hat{x}=\sum_{j}j\ket{j}\bra{j}$ | and an initial state $\hat{\rho}_{\mathrm{i}}=\ket{1}\bra{1}$ | as a function of the number $N$ of feedback cycles. As shown in the figure, the chiral transport due to topological feedback control is robust against disoder as long as $|W_{i}|\ll V$, whereas the particle transport in the non-topological case is quickly suppressed due to the Anderson localization [91] in the presence of disorder. Thus, topological feedback control can enhance the stability of particle transport against disorder.
Figure 10: Expectation value of the position as a function of the number of feedback cycles. Red (black) dots show the result for topological (non-topological) feedback control. The parameters are set to be $L=20$, $J=1$, $\tau=1$, and $V=10$. The disorder potential $W_{i}$ at each site is randomly sampled from a uniform distribution in $[-W,W]$. We take $W=0$ in panel (a) and $W=1$ in panel (b). The results are averaged over 100 samples of disorder realization.
@@ V Symmetry classification of feedback control
Symmetry enriches topological phases since two topological phases may be distiguished in the presence of symmetry [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we discuss symmetry of quantum channels for discrete quantum feedback control. In Sec. V.1, we discuss unitary symmetry. In Sec. V.2, we consider antiunitary symmetry associated with every quantum channel. While the contents in these two subsections can be found in literature [92, 93, 33, 94, 66, 67], we reproduce them to make the present paper self-contained. In Sec. V.3, we discuss the Bernard-LeClair (BL) symmetry classes of quantum channels for discrete quantum feedback control. In particular, we find that only a subset of the BL symmetry classes are compatible with projective measurement. Consequently, we obtain the ten-fold symmetry classification of discrete quantum feedback control, as listed in Table 2.
@@ V.1 Unitary symmetry
We first consider unitary symmetry of quantum channels. Similarly to the case of open quantum systems described by the Lindblad equation [92, 93], a unitary symmetry of a quantum channel can be either weak symmetry or strong symmetry [94]. A CPTP map $\mathcal{E}$ has unitary symmetry if it commutes with a unitary superoperator $\mathcal{V}$, that is,
@@
Suppose that every Kraus operator of a CPTP map commutes with a unitary operator $\hat{V}$ as $[\hat{K}_{m},\hat{V}]=0\ (\forall m)$. Then, the CPTP map satisfies Eq. (57) with $\tilde{\mathcal{V}}=\hat{V}\otimes\hat{I}$ or $\tilde{\mathcal{V}}=\hat{I}\otimes\hat{V}^{*}$. Such symmetry is called strong symmetry [92, 93, 94], which gives a more stringent condition than weak symmetry. Strong symmetry is related to a conserved quantity of the dynamics. To see this, we consider the case of continuous symmetry and assume that every Kraus operator $\hat{K}_{m}$ commutes with $\hat{V}_{\theta}=\exp[-i\theta\hat{A}]$ for all $\theta\in\mathbb{R}$, where $\hat{A}$ is a Hermitian operator. This assumption implies $[\hat{K}_{m},\hat{A}]=0\ (\forall m)$. Then, for an arbitrary density matrix $\hat{\rho}$, we have
@@ V.2 Modular conjugation symmetry
The modular conjugation operator $\tilde{J}$ is a combination of a swap operation on states and the complex conjugation of their coefficients. The modular conjugation satisfies $\tilde{J}^{2}=1$ and acts on the vectorized density matrix [see Eq. (22)] as
@@ V.3 Ten-fold symmetry classification of feedback control with projective measurement
The topological classification of non-Hermitian matrices is performed on the basis of the Bernard-LeClair (BL) symmetry classes [75], which are the non-Hermitian generalization [34, 35] of the Altland-Zirnbauer (AZ) symmetry classes [96]. Here, we discuss the BL symmetry classes of quantum channels for discrete quantum feedback control. In general, a symmetry of a CPTP map $\mathcal{E}$ that belongs to a BL symmetry class has one of the following representations [68]:
@@
where $\hat{P}_{m}$ is a projection operator satisfying $\hat{P}_{m}^{\dagger}=\hat{P}_{m}$ and $\hat{P}_{m}\hat{P}_{m^{\prime}}=\delta_{m,m^{\prime}}\hat{P}_{m}$. A key property of the projective measurement channel (64) is that its eigenvalues are either zero or one, which follows from the Hermiticity and idempotency of projection operators (see Appendix C). Physically, this property is related to the repeatability of projective measurements, that is, repeated projective measurements after the first projective measurement with outcome $m$ always yield the same outcome $m$ [47]. To identify the possible symmetry classes of the projective measurement channel (64), we note that a certain BL symmetry places a constraint on the eigenvalues of a CPTP map. Let $\xi$ be an eigenvalue of a CPTP map $\mathcal{E}$. Then, if this CPTP map has $P$ symmetry (63a) or $C$ symmetry with $\epsilon_{C}=-1$, $-\xi$ is also an eigenvalue of $\mathcal{E}$ [68]. Similarly, if a CPTP map $\mathcal{E}$ has $K$ ($Q$) symmetry, $\epsilon_{K}\xi^{*}$ ($\epsilon_{Q}\xi^{*}$) is also an eigenvalue of $\mathcal{E}$. Since a projective measurement channel (64) cannot have an eigenvalue $-1$, neither $P$ symmetry nor $X=C,K,Q$ symmetry with $\epsilon_{X}=-1$ can be a symmetry of the projective measurement channel (see Fig. 11). Thus, the possible symmetries of a CPTP map for projective measurement are the following ones:
@@
Figure 11: Eigenspectrum of a projective measurement channel (shown by green dots) consisting of eigenvalues zero and one. Since a projective measurement channel cannot have eigenvalue $-1$, its possible BL symmetry classes are restricted to the ten symmetry classes listed in Table 2 which constitute a subset of the entire 38 symmetry classes.
Now let us consider quantum channels for discrete quantum feedback control with projective measurement. Specifically, we consider a CPTP map of the form
@@
We list the ten-fold symmetry classification of quantum channels in Table 2, where each symmetry class is labeled according to the general classification of non-Hermitian symmetry classes in Ref. [34]. We note that these ten symmetry classes are equivalent to the AZ classes in Ref. [34] if we multiply $i$ to the original CPTP maps. In fact, the $C$ symmetry (65a) and the $K$ symmetry (65b) are equivalent to “time-reversal symmetry” and “particle-hole symmetry” of $i\tilde{\mathcal{E}}$, respectively [97, 34]. In Table 2, we also show the AZ classes corresponding to the symmetry classes of CPTP maps.
Once the relevant symmetry classes are identified, we can exploit the topological classification of general non-Hermitian matrices [34, 35] to examine whether there exists topological feedback control that is described by a topologically nontrivial Bloch matrix in a given symmetry class and spatial dimension. We note, however, that the Bloch matrix cannot be a general non-Hermitian matrix due to the completely positive and trace-preserving nature of quantum channels (see Appendix A.4). In particular, the complete positivity may place nontrivial constraints on possible topological phases realized with feedback control. We will present an example of this phenomenon in Sec. VI.2.
The present symmetry classification applies to the case with multipoint projective measurements. For example, let us consider the following CPTP map for feedback control with a two-point projective measurement:
@@
Even if a measurement is not projective, other symmetry classes are still not allowed unless eigenvalues of a CPTP map for measurement always appear as pairs like $(\xi,-\xi)$ or $(\xi,-\xi^{*})$ [see the paragraph above Eqs. (65a)-(65c)]. For example, if a measurement is not projective due to an error in the measurement process, the possible symmetry classes are not altered because we can take an error-free limit under which the symmetry is intact.
The ten AZ symmetry classes are formally equivalent to the symmetry classes for generators of Markovian quantum dynamics discussed in Ref. [65]. In fact, the symmetry classes that are given in Table 2 agree with the symmetry classes of CPTP maps that are generated by Lindbladians (see Appendix D). However, we note that a CPTP map for feedback control cannot have a generator if it has the zero eigenvalue [see Eq. (39)]. In general, a CPTP map generated by a Lindbladian can be classified by the 38 BL symmetry classes if one includes shifted symmetries [98, 66, 67]. However, a projective measurement channel (64) cannot have those shifted symmetries because the number of zero eigenvalues is different from that of unit eigenvalues, while shifted symmetry is accompanied by the symmetry of the eigenspectrum with respect to a certain line [98, 66, 67]. Thus, we conclude that the ten-fold symmetry classification exhausts all relevant symmetry classes of discrete quantum feedback control with projective measurement.
We note that a CPTP map always has the modular conjugation symmetry [see Eq. (62) in Sec. V.2]. The modular conjugation symmetry corresponds to a $K$ symmetry with $\eta_{K}=+1$. Thus, a quantum channel as a whole always has this symmetry. However, if a quantum channel possesses an additional unitary symmetry, special care must be taken. This is because the quantum channel can then be block-diagonalized according to the symmetry eigenvalues, and each symmetry sector (i.e. each block) does not necessarily have to have the modular conjugation symmetry. In this case, we must consider an appropriate BL symmetry class in each sector. We will revisit this point when we discuss concrete examples in Sec. VI.
Here, we provide sufficient conditions for the BL symmetries of CPTP maps to satisfy Eqs. (65a)-(65c), which are helpful to gain an insight into feedback control with symmetry. The symmetries in Eqs. (65a), (65b), and (65c) are satisfied if Kraus operators satisfy